P15-4A

r mortgage note on Decem- P154A Talkington Electronics issues a $400,000, 8%, 15-yea schedule and journal entriesber 31, 2016. The proceeds from the note are to be used in financing a new research for a mortgage note rato Prepare instal ent payments (LO 3,4) b) December 31 debit ble. ratory. The terms of the note provide for annual installment payments, exclusive of re estate taxes and insurance, of $59,612. Payments are due on December 31. Instructions (a) Prepare an installment payments schedule for the first 4 years. Mortgage Payable $27,612 (b) Prepare the entries for (1) the loan and (2) the first installment payment.
Problems: Set A 681 (c) Show how the total mortgage liability should be reported on the balance sheet at () Current liability-2017 December 31, 2017. $29,821
Answer #1

A) Bond Discount Amortization

Annual Interest Periods

Interest to be Paid

Interest expense

Discount Amortization

Principal balance

(A)

(B) =(D×8%)

C = (B - A)

D = (D - C)

Issue date

$400,000

1

$59,612

$32,000

$27,612

$372,388

2

$59,612

$29,791

$29,821

$342,567

3

$59,612

$27,405

$32,207

$310,360

4

$59,612

$24,829

$34,783

$275,577

B)

Particulars

Debit

Credit

1)

Cash

400,000

Mortgage payable

400,000

2)

Interest expenses

32,000

Mortgage payable

27,612

Cash

59,612

 

Computation of interest expenses on mortgage payable:

Interest expenses = Mortgage loan payable * Interest rate * Time period

=$400,000 * 8% * 1

=$32,000

Thus, interest expenses on mortgage payable equals $32,000.

Computation of mortgage payable:

Mortgage payable =Cash payment - Interest expenses

=$59,612 - $32,000

= $27,612

Thus, mortgage payable equals $27,612

3) Current liability is categorised to be obligation that the an organisation is required to pay from the remaining current assets or creation of other current liabilities within the operating cycle or a fiscal year whichever is higher. Long-term liabilities are obligations that the organisation is required to pay after at least one year or more. These are otherwise termed to be the long-term debt

Balance Sheet (Partial)

Particulars

Amount

Amount

Liabilities:

Current liabilities:

Current portion of mortgage payable

29,821

Total current liabilities

29,821

Long-term liabilities:

Mortgage payable

342,567

Total long-term liabilities

342,567

Total liabilities

342,567